# Missing Data Imputation using Neural Cellular Automata

### v1

@@ 1 Introduction
Our contributions are: (i) We introduce NICA, a novel generative imputation model based on NCA architecture; (ii) We provide several experiments to compare the performance of our model with other imputing algorithms at different settings; (iii) We illustrate the clustering behavior of the data samples throughout the growing process; (iv) We report accuracy in downstream task when training with our imputed data versus other methods.
@@ 4.1 Model
More particularly, similar to other imputation methods, NICA receives as input a matrix $\tilde{\mathbf{X}}$, which contains blank cells. First, it creates a matrix $\mathbf{M}$ of the same size as $\tilde{\mathbf{X}}$ to mark the observed positions. Next, the data is normalized and pre-assigned by replacing missing values with 0. After that, the initial seed is constructed by randomly corrupting the pre-assigned data at the known cells, for a certain percentage. Now, we have a corrupted data $\mathbf{X}_{c}$ and its corresponding corrupting mask $\mathbf{M}_{c}$.
@@
After the loop, grown data $\mathbf{X}_{c}^{(K)}$ is used to calculate losses. Our objectives are (i) recovering missing data from the corrupted data, (ii) generating values that are as close as observed data as possible, and (iii) staying away from trivial solution. For the first purpose, we calculate the recovering loss
$\mathcal{L}_{recovering}=\mathbb{E}_{\mathbf{M}_{c}}\left[\mathcal{L}\left(\mathbf{X}_{c}^{(K)},\mathbf{X}_{c}\right)\right],$
where $\mathcal{L}\left(\mathbf{X},\mathbf{Y}\right)$ is the squared difference between two matrices $\mathbf{X}$ and $\mathbf{Y}$. This loss is partly adopted from the objective in training NCA architecture. Then, a dominant loss in missing data imputation literature is the difference between generated data and known data, which we call observed loss
$\mathcal{L}_{observed}=\mathbb{E}_{\mathbf{M}}\left[\mathcal{L}\left(\mathbf{X}_{c}^{(K)},\mathbf{X}_{c}\right)\right].$
@@
$\mathcal{L}_{missing}=-\mathbb{E}_{\mathbf{1-M}}\left[\mathcal{L}\left(\mathbf{X}_{c}^{(K)},\mathbf{X}_{c}\right)\right].$
@@
$\mathcal{L}_{model}=\eta_{1}\mathcal{L}_{recovering}+\eta_{2}\mathcal{L}_{observed}+\mathcal{L}_{missing},$
where $\eta_{1}$ and $\eta_{2}$ are the hyperparameters to adjust the contribution of $\mathcal{L}_{recover}$ and $\mathcal{L}_{observed}$ in comparison with $\mathcal{L}_{missing}$. When $\mathcal{L}_{model}$ has been optimized, the generated data is $\bar{\mathbf{X}}$, and the assigned data is calculated by
$\hat{\mathbf{X}}=\mathbf{M}\odot\tilde{\mathbf{X}}+(\mathbf{1-M})\odot\bar{\mathbf{X}},$
@@ 5.1 Experimental Design
In our experiments, missing data is standardized and $20\%$ of the observed part is randomly corrupted for $v=8$ times. We train each data $1000$ iterations with the batch size of $1024$, and the number of growing steps is $K=10$. Since the $\operatorname{\texttt{CosineAttention}}$ outputs a tensor with the same shape of its input, output of $\operatorname{\texttt{stack}}$ is double the original dimension. The $\operatorname{\texttt{FNN}}$ is a fully-connected feedforward network with $2$ linear layers and ReLU activation. The first linear layer maps the state information to the hidden space, whose dimension if $5$ times higher than that of the original, then the later one projects its back to original space. At each growing step, $10\%$ of samples is skipped from updating. When calculating model loss, we choose $\eta_{1}=\eta_{2}=10$ to emphasize the importance of $\mathcal{L}_{recovering}$ and $\mathcal{L}_{observed}$ in imputation task. We train our model by optimizing $\mathcal{L}_{model}$ using Adam algorithm [11] with learning rate $0.001$.

### v2

@@ 1 Introduction
Our contributions are: (i) We introduce NICA, a novel generative imputation model based on NCA architecture; (ii) We provide several experiments to compare the performance of our model with other imputing algorithms at different settings; (iii) We illustrate the clustering behavior of the data samples throughout the growing process; (iv) We report accuracy in downstream task when training with our imputed data versus other methods. To facilitate reproducibility, we provide the source code at: github.com/TrungTin98/NICA.
@@ 4.1 Model
More particularly, similar to other imputation methods, NICA receives as input a matrix $\tilde{\mathbf{X}}$, which contains blank cells. First, it creates a matrix $\mathbf{M}$ of the same size as $\tilde{\mathbf{X}}$ to mark the observed positions. Next, the data is normalized and pre-assigned with zeros to create a simple baseline version denoted by $\bar{\mathbf{X}}_{0}$. After that, the initial seed is constructed by randomly corrupting a certain proportion of the known cells of $\bar{\mathbf{X}}_{0}$. Now, we have a corrupted data $\mathbf{X}_{c}$ and its corresponding corrupting mask $\mathbf{M}_{c}$.
@@
After the loop, grown data $\mathbf{X}_{c}^{(K)}$ is used to calculate losses. Our objectives are (i) recovering missing data from the corrupted data, (ii) generating values that are as close as observed data as possible, and (iii) staying away from trivial solution. For the first purpose, we calculate the recovering loss, which is the difference between the grown data and our zero-imputed version at the corrupted cells
$\mathcal{L}_{recovering}=\mathbb{E}_{\mathbf{M}_{c}}\left[\mathcal{L}\left(\mathbf{X}_{c}^{(K)},\bar{\mathbf{X}}_{0}\right)\right],$
where $\mathcal{L}\left(\mathbf{X},\mathbf{Y}\right)$ is the squared difference between two matrices $\mathbf{X}$ and $\mathbf{Y}$. This loss is partly adopted from the objective in training NCA architecture. Then, a dominant loss in missing data imputation literature is the dissimilarity between generated data and known data, which we call observed loss
$\mathcal{L}_{observed}=\mathbb{E}_{\mathbf{M}}\left[\mathcal{L}\left(\mathbf{X}_{c}^{(K)},\bar{\mathbf{X}}_{0}\right)\right].$
@@
$\mathcal{L}_{missing}=-\mathbb{E}_{\mathbf{1-M}}\left[\mathcal{L}\left(\mathbf{X}_{c}^{(K)},\bar{\mathbf{X}}_{0}\right)\right].$
@@
$\mathcal{L}_{model}=\alpha_{1}\mathcal{L}_{recovering}+\alpha_{2}\mathcal{L}_{observed}+\mathcal{L}_{missing},$
where $\alpha_{1}$ and $\alpha_{2}$ are the hyperparameters to adjust the contribution of $\mathcal{L}_{recover}$ and $\mathcal{L}_{observed}$ in comparison with $\mathcal{L}_{missing}$. When $\mathcal{L}_{model}$ has been optimized, the generated data is $\bar{\mathbf{X}}$, and the assigned data is calculated by
$\hat{\mathbf{X}}=\mathbf{M}\odot\bar{\mathbf{X}}_{0}+(\mathbf{1-M})\odot\bar{\mathbf{X}},$
@@ 5.1 Experimental Design
In our experiments, missing data is standardized and $20\%$ of the observed part is randomly corrupted for $v=8$ times. We train each data $1000$ iterations with the batch size of $1024$, and the number of growing steps is $K=10$. Since the $\operatorname{\texttt{CosineAttention}}$ outputs a tensor with the same shape of its input, output of $\operatorname{\texttt{stack}}$ is double the original dimension. The $\operatorname{\texttt{FNN}}$ is a fully-connected feedforward network comprising $2$ linear layers with ReLU activation and a dropout rate of $0.5$. The first linear layer maps the state information to the hidden space, whose dimension if $5$ times higher than that of the original, then the later one projects its back to original space. At each growing step, $10\%$ of samples is skipped from updating. When calculating model loss, we choose $\alpha_{1}=\alpha_{2}=10$ to emphasize the importance of $\mathcal{L}_{recovering}$ and $\mathcal{L}_{observed}$ in imputation task. We train our model by optimizing $\mathcal{L}_{model}$ using Adam algorithm [11] with learning rate $0.001$.